slove the equation 4y+228=352



Answer by usyim88hk(158)

4y+228=352
First of all, subtract 228 from both sides
4y + 228 - 228 = 352 - 228
4y = 124
Then divide both sides by 4
4y/4 = 124/4
y = 31